My son can clean the attic in 4 hours and I can do it in 2 hours. How long would it take if we worked together.
Answer by nerdybill(7384)

My son can clean the attic in 4 hours and I can do it in 2 hours. How long would it take if we worked together.
.
Let x = time (hours) for both
then
x(1/4 + 1/2) = 1
x(1/4 + 2/4) = 1
x(3/4) = 1
x(3) = 4
x = 4/3 hours
or
x = 1 hour and 20 minutes